# Safe Control using Occupancy Grid Map-based Control Barrier Function (OGM-CBF)

### v2

@@ V-A CARLA
The experiment aims to validate the safe navigation of an autonomous car using OGM-CBF in a scenario including diverse obstacles. The control inputs are $\boldsymbol{u}=[\mathrm{v},\boldsymbol{\omega}]^{T}$, and steering and throttle commands are calculated using the standard kinematic bicycle model in a 2D workspace [28]. A LiDAR sensor is mounted on the ego vehicle to generate the occupancy grid map $m$ online, where each grid cell represents a $0.1$ by $0.1$ meter area of the environment. The simulation involves navigating an unknown environment cluttered with various stationary obstacles, such as cars and pedestrians.
In this experiment, safe navigation is deployed by a controller based on Algorithm (1), where a quadratic Lyapunov function is used to align the vehicle’s heading with a target heading, while the optimization’s cost function maintains a desired speed $\mathrm{v}_{d}=2m/s$. The controller’s effectiveness can be inspected from Fig. 2. In this Figure, in the first row, the orange line shows the path of the ego vehicle. In addition to cars, smaller obstacles such as pedestrians, bicyclists, and motorcyclists are spawned in the environment (marked with red circles). The second row illustrates the upper half of the ego-centric local occupancy grid map $m$ during the experiment, and the last row displays $\Phi_{s}(\boldsymbol{x},m)$ and the gradient field of it in the map frame for all the states in the map. It should be noted that in Fig. 2, the ego vehicle initiates its journey from the right side of the page and proceeds towards the left. Additionally, as the robot has a positive lower bound velocity, $T(\phi(\boldsymbol{x},M_{b}(m)))$ is only constructed for the upper half of the ego-centric map as the function $p(a\tanh(b\phi(\boldsymbol{x},M_{b}(m))))$, where $a$ and $b$ are tuned based on the desired behaviour, and $p(.)$ is a polynomial interpolation function as described in section IV.
Fig. 3 quantitatively illustrates that the designed controller in fact navigates safely during the experiment. The first row displays the value of $h(\boldsymbol{x})$, and the second row shows $\dot{h}(\boldsymbol{x},\boldsymbol{u})+\alpha(h(\boldsymbol{x}))$, which remains non-negative, indicating constant satisfaction of the safety constraint. The third and fourth rows depict the control inputs $\boldsymbol{u}=[\mathrm{v},\boldsymbol{\omega}]^{T}$ generated by Algorithm (1) (solid lines), alongside the actual linear and angular velocities of the vehicle (dashed lines) with respect to time. Note the value of $h(\boldsymbol{x})$ can become negative near obstacles due to the approximation of the system (1) as a kinematic model.
@@ V-B Real-world
Besides simulated validation, additional experiments were conducted on an industrial mobile robot (MiR100), which is a differential wheeled robot equipped with an Nvidia Jetson AGX Orin and two laser scanners. Mapping and safe control are deployed using the ROS2 framework on this platform, and localization is done internally using onboard sensors.
In this experiment, the robot utilizes a pre-established global map of the environment, which is dynamically updated with an average frequency of $16.372$ Hz using real-time sensor measurements collected from laser scanners. Each grid cell in the map covers a $0.05$ by $0.05$ meter section of the environment. Safe navigation is managed by a controller based on Algorithm (1) using OGM-CBF, with a quadratic Lyapunov function for maintaining the desired heading. Moreover, the optimization’s cost function is designed to keep the speed of the robot at $\mathrm{v}_{d}=0.25$ m/s. The employed controller runs on the platform with an average frequency of $12.498$ Hz, and the applied function $T(\phi(\boldsymbol{x},M_{b}(m)))$ is the same as in the CARLA experiments.
@@
Fig. 4 shows the robot’s path (orange lines) during the experiment. The left image demonstrates OGM-CBF’s capabilities to avoid obstacles with arbitrary shapes (the blue obstacle and a human). The middle image shows the global occupancy grid map updated in real-time by laser scanners, and the robot’s path in the map frame. The right image presents $\Phi_{s}(\boldsymbol{x},m)$ and its gradient field in the map frame for all the states in the map (blue arrows). Moreover, the gradients of $\Phi_{s}(\boldsymbol{x},m)$ along the robot’s path are marked by red arrows, indicating the direction of avoidance near obstacles utilized in the controller. Note that the robot progresses from the top to the bottom of the images throughout the experiment.
Fig. 5 quantitatively validates the robot’s safe navigation using OGM-CBF during the experiment. The first row displays the value of $h(\boldsymbol{x})$, the second row presents $\dot{h}(\boldsymbol{x},\boldsymbol{u})+\alpha(h(\boldsymbol{x}))$, and the final two rows illustrate the commanded (solid blue line) and actual (dashed red line) linear and angular velocities of the robot with respect to time.
@@ VI CONCLUSION
This paper introduces a novel method, the Occupancy Grid Map-based Control Barrier Function (OGM-CBF), which enhances safe navigation in robotics across both known and unknown terrains. By integrating the Control Barrier Function (CBF) with occupancy grid maps, our approach effectively accommodates arbitrary obstacle shapes and configurations. The designed CBF is sensor-agnostic and can be implemented using a range of hybrid sensor modalities. Moreover, the OGM-CBF is formulated as a single constraint in a Quadratic Programming (QP) formulation, taking into account all previously seen obstacles. Additionally, the constructed CBF considers the robot’s size while applying established inflation methods to the occupancy grid map.
We validated the effectiveness of OGM-CBF through experiments in the CARLA simulator and practical implementations on an industrial mobile robot, confirming its capability for reliable autonomous navigation. Future work will demonstrate the integration of different sensor modalities where safe/unsafe segmentation is integrated to one OGM, and incorporate the effect of the actuator dynamics into the CBF design.

### v3

@@ V-A CARLA
The experiment aims to validate the safe navigation of an autonomous car using OGM-CBF in a scenario including diverse obstacles. In this case, the state of the system is $\boldsymbol{x}=[x,y,\psi]^{T}$ (the coordinates of the vehicle’s center in the 2D plane, and the heading of the vehicle), and the workspace of the system is two dimensional, with $\boldsymbol{u}=[v,\omega]^{T}$ (linear and angular speeds) as the control inputs. Steering and throttle commands are calculated using the standard kinematic bicycle model in a 2D workspace [27]. A LiDAR sensor is mounted on the ego vehicle to generate the occupancy grid map $m$ online, where each grid cell represents a $0.1$ by $0.1$ meter area of the environment. The LiDAR sensor’s range is 70 meters, with $760000$ points per scan and $35$ degrees field of view, and the map binarization threshold $\mu=0.5$. The simulation involves navigating through an unknown environment cluttered with various stationary obstacles, such as cars and pedestrians.
Our control objective is to maintain the vehicle heading and avoid obstacles. Safe navigation is deployed by a controller based on Algorithm (1) and the optimization framework (14). The performance controller is defined by an output tracking error as a Lyapunov function $V(\boldsymbol{x})=\frac{1}{2}(\psi-\psi_{ref})^{2}$, where $\psi$ is the robot heading (angle of $\hat{x}$), while the optimization’s cost function is $J(\boldsymbol{u},\boldsymbol{x},\delta)=\frac{1}{2}(v-v_{ref})^{2}+\frac{1}{2}(\omega-\omega_{ref})^{2}+\delta^{2}$, where reference velocities are $\mathrm{v}_{ref}=2m/s$, $\mathrm{\omega}_{ref}=-k(\psi-\psi_{ref})$. The controller’s effectiveness can be inspected from Fig. 2. In this Figure, in the first row, the orange line shows the path of the ego vehicle. In addition to cars, smaller obstacles such as pedestrians, bicyclists, and motorcyclists are spawned in the environment (marked with red circles). During the experiment, OGM-CBF utilizes an upper half of the ego-centric local occupancy grid map as $m$, shown in the second row. And the last row displays $\Phi_{s}$ and the gradient field of it in the map frame.
Additionally, as the robot has a positive lower bound velocity, $T(\phi(\boldsymbol{p},\mathbb{X}_{s}))$ is only constructed for the upper half of the ego-centric map as the function $c(a\tanh(b\phi(\boldsymbol{p},\mathbb{X}_{s}))))$, where $a=3$ and $b=0.01$ are tuned based on the desired behaviour, and $c(.)$ is a polynomial interpolation function as described in section III-A. Moreover, $l_{s}=-0.5$ and $l_{a}=0.5$. Fig. 3 quantitatively illustrates that the designed controller in fact navigates safely during the experiment. The first row displays the value of $h(\boldsymbol{x})$, and the second row shows $\dot{h}(\boldsymbol{x},\boldsymbol{u})+\alpha(h(\boldsymbol{x}))$, which remains non-negative, indicating constant satisfaction of the safety constraint. The third and fourth rows depict the control inputs $\boldsymbol{u}=[\mathrm{v},\omega]^{T}$ generated by Algorithm (1) (solid lines), alongside the actual linear and angular velocities of the vehicle (dashed lines) with respect to time. Note the value of $h(\boldsymbol{x})$ can become negative near obstacles due to the approximation of the system (1) as a kinematic model.
@@ V-B Real-world
Besides simulated validation, additional experiments were conducted on an industrial mobile robot (MiR100), which is a differential wheeled robot equipped with an Nvidia Jetson AGX Orin and two onboard compact safety laser scanner SICK S300, with scanning angle of $270$ degrees. The system is considered with 2D workspace, same as Section V-A. Mapping and safe control are deployed using the ROS2 framework on this platform, and localization is done internally using onboard sensors.
In this experiment, the robot utilizes a pre-established global map of the environment, which is dynamically updated with an average frequency of $16.372$ Hz using real-time sensor measurements collected from laser scanners. Each grid cell in the map covers a $0.05$ by $0.05$ meter section of the environment with map binarization threshold of $\mu=0.5$. Safe navigation is managed by a controller based on Algorithm (1) using OGM-CBF and the optimization framework (14), with the same quadratic Lyapunov function and cost function as in Section V-A, which is similarly utilized to maintain the desired heading and $\mathrm{v}_{ref}=0.25$. The employed controller runs on the platform with an average frequency of $12.498$ Hz, and the applied function $T(\phi(\boldsymbol{p},\mathbb{X}_{s}))$ is the same as in the CARLA experiments, with $a=3$ and $b=0.01$. Moreover, $l_{s}=-0.35$ and $l_{a}=0.35$.
@@
Fig. 4 shows the robot’s path (orange lines) during the experiment. The subfigure (a) demonstrates OGM-CBF’s capabilities to avoid obstacles with arbitrary shapes (the black obstacle and a human). (b) shows the global occupancy grid map updated in real-time by laser scanners, and the robot’s path in the map frame. (c) presents $\Phi_{s}$ and its gradient field in the map frame (black arrows). Moreover, the gradients of $\Phi_{s}$ along the robot’s path are marked by red arrows, indicating the direction of avoidance near obstacles utilized in the controller.
Fig. 5 quantitatively validates the robot’s safe navigation using OGM-CBF during the experiment. The first row displays the value of $h(\boldsymbol{x})$, the second row presents $\dot{h}(\boldsymbol{x},\boldsymbol{u})+\alpha(h(\boldsymbol{x}))$, and the final two rows illustrate the commanded (solid black line) and actual (dashed red line) linear and angular velocities of the robot with respect to time.
@@ VI CONCLUSION
This paper introduces OGM-CBF, a novel method for safe robot navigation in both known and unknown environments. By combining Control Barrier Functions (CBF) with Occupancy Grid Maps (OGM) and Signed Distance Functions (SDF), the method efficiently handles arbitrary obstacle shapes and sensor inputs. OGM-CBF is implemented as a single constraint in a Quadratic Programming (QP) formulation. Validation in the CARLA simulator and on an industrial mobile robot demonstrates reliable autonomous navigation. Future work will focus on considering actuator dynamics in the CBF design, and including a method for parameter tuning given certain desired robot behaviour in the presence of obstacles.